# 00footnotetext: Shared first author; Shared last author

### v3

@@ Appendix I Model-based quality filters
Appendix N Decontamination
In Section 4.6, we examined the results of the contamination analysis with respect to MMLU. Here, we present some analysis with respect to the other validation sets. Instead of the MMLU-specific decontamination that we performed in Section 4.6, here we follow a more general approach based on token overlaps.
Overall, a generally applicable decontamination rule is difficult to specify, given the potentially subjective nature of what constitutes as contamination in text data as well as the diversity in formats across tasks. Following Touvron et al. [162], we search for contaminated tokens that exist in overlapping 10-grams (or longer) between DCLM-baseline and our downstream tasks. We measure the percentage of samples in each evaluation set where more than 80% of the tokens are contaminated (such samples are considered “dirty” per Touvron et al. [162]), as well as the percentage where less than 20% of the tokens are contaminated (considered “clean” by the same criterion).
We examine the difference of performance of the same 7B-2x model trained on DCLM-baseline, between the full evaluation set and the evaluation samples that are marked as “not dirty” per the criterion in Touvron et al. [162] (less than 80 % of the tokens in the sample are marked as contaminated), and between the full evaluation set and samples marked as “clean” using the same criterion (less than 20 % of the tokens are marked). Results can be seen in Figure 10, where we see that the difference in performance over the full dataset and the “not dirty” samples is minimal. In fact, for BoolQ and SQuAD, which are marked as highly contaminated, our model performs slightly better on the “not dirty” subset. Moreover, in Figure 11 we see that the difference in performance between the full evaluation set and the “clean” subset is similarly small for most datasets. We note here that it’s difficult to identify a correct threshold for what counts as a contaminated sample (as 20 % token overlap might lead to many false positives, but at the same time 80 % might be too high to detect all contaminated samples).
Appendix O Instruction tuning
@@
Appendix P Scaling up
The final run 2.5T shown in Figure 1 and Table 9 uses a two stage training procedure as followed in Groeneveld et al. [68], Hu et al. [74] and Team et al. [154]. For stage 1 we use the hyper-parameters from Table 27.
@@
After 2T tokens, we cooldown on a re-weighted pre-training distribution. For the cooldown distribution we use a mix of 70% DCLM-baseline with a tighter fastText threshold (top 7% rather than top 10%) and 30% ProofPile. We keep all the hyperparameters the same as Table 27, so we cooldown to the same final learning rate, just over a smaller number of tokens. Before the cool-down MMLU performance was approximately 52%, and the LR was approximately $1\times 10^{-3}$.
We performed 2 independent cooldowns, one for 270B tokens and another for 200B tokens, and created a “model soup" [173] with a weight of 0.8 on the 270B cooldown and a weight of 0.2 on the 200B cooldown. Thus we the total number of tokens seen by this model is 2.5T. We present results of each individual cooldown and the model soup in Table 28. The model in Figure 1 and Table 9 uses the final model soup after long-context training for 100B tokens as described in Section P.2.
@@
In Appendix O we show how instruction tuning the above “model soup" for 80B additional tokens leads to strong performance on instruction tuning benchmarks and out-performs instruction tuned variants of similar 7B models such as Gemma-7B. In addition to the IT benchmarks covered in Appendix O, In Table 29 we show that a small amount of instruction tuning provides large improvements in “Extended" evals at the cost of a small degradation in “Core" and “MMLU" evals. Notably we note that our GSM8k performance goes 2.5% to 52.5% which is comparable to other similar language models that mixed IT data into pretraining such as Gemma-7B.
@@
We show that the above strategy results in similar performance on regular evaluations as the starting checkpoint and significantly improves on the multi-document question-answering evaluation. We use the evaluation setup described in [95]: the context is filled with $k$ documents followed by a question. We ensure that one of the $k$ documents includes the answer to the question (a.k.a., golden document). We use $k=1,10,20,30$, and for each case, we run the evaluation multiple times by changing the position of the golden document in the context and report the average. Results are reported in table 30. We demonstrate that long context adaptation results in a checkpoint (DCLM-8k) that matches the original model on regular evaluations and significantly improves multi-document QA showing its long-context capabilities.
@@
Appendix Q Account of compute costs
@@
We note the compute cost of training runs for each competition scale in Table 1. In total, we estimate that our runs for DCLM sum up to approximately 1M H100 hours. Our precise estimate from our experimental test bed is 772K H100 hours for training, but this is likely an underestimate due to additional compute that was not tracked, such as due to training failures.
Appendix R Existing assets used
@@
In addition to the above, as described in Sections 5, 4.5 and 4.4 and Appendices L and O we make use of the following datasets:
@@
    RedPajama [160] is used for ablations in Section 4.5 and Appendix L. Note that we do not release models or datasets that include this data. RedPajama filters the datasets it uses keeping only permissive licenses, and refers the user to adhere to underlying licenses where appropriate, as described in https://huggingface.co/datasets/togethercomputer/RedPajama-Data-1T.
@@
    • The purpose of DCLM and the associated DCLM-Pool and DCLM-baseline datasets is to enable the study of what makes a strong pretraining dataset for large language models. These models are transformative to society and act as the foundation of numerous applications, but they are often associated with steep costs. While prior work explores many curation techniques, it is often coupled with various architectural and training design choices and evaluated in different settings, making controlled comparison nearly impossible. This slows down progress and forces a lot of duplicate work between research teams. Prior work mainly focuses on data curation in the context of supervised datasets and smaller scales (see Section 2 and Appendix B). In our initial release of DCLM, we focus on 53 downstream language understanding tasks that also include reasoning abilities, math, code, and more. For details see Section 3.5 and Appendix G.
@@
    • DCLM-Pool and DCLM-baseline were created by a group of researchers with the following affiliations, listed in alphabetical order: Allen Institute for Artificial Intelligence, Apple, Carnegie Mellon University, Columbia University, Contextual AI, Cornell University, DatologyAI, Harvard University, Hebrew University, Juelich Supercomputing Center, Research Center Juelich, SambaNova Systems, Stanford University, SynthLabs, Tel Aviv University, Toyota Research Institute, TU Munich, University of California, Los Angeles, University of California, Santa Barbara, University of Southern California, The University of Texas at Austin, University of Washington.
@@
    • We anticipate that DCLM benchmark, tooling and pools will drive data-centric research in ML and AI, fostering the development of the next generation of web-scale datasets, enhancing model abilities, lowering training costs and develop knowledge sharing across research teams.
@@
    • Each instance represented a web-crawled page (document). It contains the URL and the corresponding HTML content. Each sample is also tagged with metadata about its crawl time and additional information such as the detected language, for processed instances such as those in DCLM-baseline. Additional information can be found in Appendix E.
@@
    • DCLM-Pool contains $\sim$200B documents, all of which are of the same instance, and comes from hundreds of millions of different sources. The subset DCLM-baseline contains approximately 3B documents.
@@
    • DCLM-Pool is an unfiltered web-text corpus comprised of all Common Crawl data prior to 2023. As such, it represent the full breadth of possible instances from this source. However, we note that Common Crawl does not cover the entire web data, due to reach and compute limitations for instance. For our DCLM-baseline, we use various filtering and deduplication strategies as described in Section 4 in the explicit attempt to improve its quality for preatining, thus removing low-quality instances, and in doing so, becoming non-representative of the full set of instances. For a complete treatment and visualization of our data processing funnel, see Sections 4, 4.3 and 4.2 and Appendix E.
@@ 8. Q12
    • No. The evaluation procedure is made of tasks as described in Section 3.5. We also attempt to prevent test set contamination in as described in Section 4.6 and Appendix N.
@@ 15. Q19
    • As names and other identifiers are frequent in web data, it is likely that some content can be linked back to specific individuals. However, in most public sites which Common Crawl scrape people publish such information willingly, knowing it will be visible and public.
@@ 16. Q20
    • Yes. DCLM-Pool is created from data that is available on the public internet. Since people often debate their political views, sexual preferences, religious beliefs and other such information, it is highly likely such information is contained in the dataset. While such information is often published willingly in the explicit intent that it will be publicly visible (see Q19), we do encourage additional research on filtering such data both to preserve privacy as well as to discard any potentially biased or toxic content from the training data of the models.
@@ 2. Q23
    • We begin by downloading the entire Common Crawl data prior to 2023. We ran Python-based processing scripts to parse these archives, filtering low-quality or irrelevant content, deduplicate samples and in some cases decontaminate against downstream tests sets, and compute various model-based features. We ran processes on hundreds of AWS CPU nodes for Common Crawl parsing and data downloading. Model-based features were run on GPU clusters. For software links see Q37 and Appendix R or refer to https://datacomp.ai/dclm.
@@ 3. Q24
    • DCLM-Pool is not a probabilistic sample. As described in Q7, DCLM-Pool contains all data from Common Crawl before 2023. Common Crawl is a sample of the Web, and we refer to Common Crawl documentation for details of their sampling process.

### v4

@@ Appendix I Architecture ablations
Appendix O Decontamination
Dataset Percentage of samples flagged
DCLM-baseline 0.007%
Dolma-V1.7 0.001%
FineWeb-Edu 0.009%
In Section 4.6, we examined whether contamination explains the strong results of DCLM-baseline on MMLU and Hellaswag. Here, we first supplement this analysis by showing in Table 25 that the percentage of documents in DCLM-baseline that are flagged by our MMLU-specific decontamination is roughly similar to that for other high-performing contemporary datasets. Next, we present some analysis with respect to other open-source datasets and evaluation tasks.
To expand our analysis to more evaluation tasks, instead of the decontamination that we performed in Section 4.6, here we follow a more general approach based on token overlaps. Overall, a generally applicable decontamination rule is difficult to specify, given the potentially subjective nature of what constitutes as contamination in text data as well as the diversity in formats across tasks. Following Touvron et al. [170], we search for contaminated tokens that exist in overlapping 10-grams (or longer) between DCLM-baseline and our downstream tasks. We measure the percentage of samples in each evaluation set where more than 80% of the tokens are contaminated (such samples are considered “dirty” per Touvron et al. [170]), as well as the percentage where less than 20% of the tokens are contaminated (considered “clean” by the same criterion).
We examine the difference in performance of the same 7B-2x model trained on DCLM-baseline between the full evaluation set and the evaluation samples that are marked as “not dirty” per the criterion in Touvron et al. [170] (less than 80 % of the tokens in the sample are marked as contaminated), and between the full evaluation set and samples marked as “clean” using the same criterion (less than 20 % of the tokens are marked). Results can be seen in Figure 11, where we see that the difference in performance over the full dataset and the “not dirty” samples is minimal. In fact, for BoolQ and SQuAD, which are marked as highly contaminated, our model performs slightly better on the “not dirty” subset. Moreover, in Figure 12 we see that the difference in performance between the full evaluation set and the “clean” subset is similarly small for most datasets. While Hellaswag appears to be a notable exception, we showed in Section 4.6 that removing contaminated Hellaswag examples from our training set does not decrease performance. We note here that it’s difficult to identify a correct threshold for what counts as a contaminated sample (as 20 % token overlap might lead to many false positives, but at the same time 80 % might be too high to detect all contaminated samples).
Appendix P Instruction tuning
@@
Appendix Q Scaling up
The final run 2.5T shown in Figure 1 and Table 8 uses a two stage training procedure as followed in Groeneveld et al. [73], Hu et al. [80] and Team et al. [162]. For stage 1 we use the hyperparameters from Table 27.
@@
After 2T tokens, we cooldown on a re-weighted pretraining distribution. For the cooldown distribution we use a mix of 70% DCLM-baseline with a tighter fastText threshold (top 7% rather than top 10%) and 30% ProofPile. We keep all the hyperparameters the same as Table 27, so we cooldown to the same final learning rate, just over a smaller number of tokens. Before the cool-down MMLU performance was approximately 52%, and the LR was approximately $1\times 10^{-3}$.
We performed 2 independent cooldowns, one for 270B tokens and another for 200B tokens, and created a “model soup" [182] with a weight of 0.8 on the 270B cooldown and a weight of 0.2 on the 200B cooldown. Thus the total number of tokens seen by this model is 2.5T. We present results of each individual cooldown and the model soup in Table 28. The model in Figure 1 and Table 8 uses the final model soup after long-context training for 100B tokens as described in Section Q.2.
@@
In Appendix P we show how instruction tuning the above “model soup" for 80B additional tokens leads to strong performance on instruction tuning benchmarks and out-performs instruction tuned variants of similar 7B models such as Gemma-7B. In addition to the IT benchmarks covered in Appendix P, In Table 29 we show that a small amount of instruction tuning provides large improvements in “Extended" evals at the cost of a small degradation in “Core" and “MMLU" evals. Notably we note that our GSM8k performance goes 2.5% to 52.5% which is comparable to other similar language models that mixed IT data into pretraining such as Gemma-7B.
@@
We show that the above strategy results in similar performance on regular evaluations as the starting checkpoint and significantly improves on the multi-document question-answering evaluation. We use the evaluation setup described in [101]: the context is filled with $k$ documents followed by a question. We ensure that one of the $k$ documents includes the answer to the question (a.k.a., golden document). We use $k=1,10,20,30$, and for each case, we run the evaluation multiple times by changing the position of the golden document in the context and report the average. Results are reported in Table 30, which show that long context adaptation results in a checkpoint (DCLM-8k) that matches the original model on regular evaluations and significantly improves multi-document QA showing its long-context capabilities.
@@
Scaling up the 1B model
In addition to scaling up our 7B model, we also train our 1.4B parameter model for 4.3T tokens to show the utility of DCLM-baseline for smaller-scale models. Similar to the 7B model, we train for 4.3T tokens on DCLM-baseline combined with the StarCoder and ProofPile2 datasets, with the hyper-parameters from Table 31. Unlike the 7B model, we train for the full 4.3T tokens; we do not perform a separate cooldown or model souping phase. We present results for this model in Table 32.
Table 31: Hyper-parameters for the 1B large-scale run.
Hyper-parameter/Config Value
Training Tokens 4,319,385,600,000
Warmup Steps 5000
Initial Learning Rate $1\times 10^{-2}$
Weight Decay 0.01
Final Learning Rate $3\times 10^{-5}$
Global Batch Size 2048
Accumulation Steps 2
Query-Key Normalization True
Z Loss $1\times 10^{-6}$
Table 32: Evaluation for our 1.4B model compared to state-of-the-art small models.
Model Params Tokens Open dataset? Core MMLU Extended
Open weights, closed datasets
Phi-1.5 1.3B 0.15T 39.8 43.4 25.7
Qwen2-1.5B 1.5B 7T 42.1 56.4 32.4
Gemma-2B 2.5B 3T 43.3 40.8 26.6
Open weights, open datasets
OLMo-1B 1.2B 3T 29.7 26.0 16.1
SmolLM 1.7B 1T 36.3 30.0 21.2
DCLM-baseline 1.4B 4.3T 45.2 47.5 28.1
Full results from Figure 1
In Table 33, we reproduce the results from Figure 1 in table format to ease future comparisons.
Table 33: We reproduce the results from Figure 1 (left) in tabular format for ease of reference below.
Model Params Tokens Open dataset? Core MMLU Extended
C4 412M 8B 13.1 24.9 8.0
Dolma v1 412M 8B 13.6 25.2 9.0
RefinedWeb 412M 8B 15.1 25.5 8.9
RedPajama 412M 8B 15.4 24.8 7.6
DCLM-Baseline (ours) 412M 8B 18.0 25.5 9.8
FineWeb-Edu 412M 8B 18.3 25.5 10.1
C4 1B 29B 23.7 25.4 12.5
RefinedWeb 1B 29B 25.1 24.0 12.9
RedPajama 1B 29B 25.7 24.3 13.5
Dolma v1 1B 29B 25.8 24.3 13.5
FineWeb-Edu 1B 29B 26.6 26.3 13.5
OLMo 1B 2T 29.7 26.0 16.1
DCLM-Baseline (ours) 1B 29B 30.2 23.8 15.4
C4 3B 56B 30.5 26.3 16.0
Dolma v1 3B 56B 31.5 27.6 16.7
RedPajama 3B 56B 31.6 25.3 16.6
RefinedWeb 3B 56B 33.2 25.2 17.3
C4 7B 138B 37.1 25.5 19.8
DCLM-Baseline (ours) 3B 56B 37.1 25.2 20.0
RedPajama 7B 138B 37.8 25.1 20.6
Dolma v1 7B 138B 38.2 25.9 19.6
FineWeb-Edu 7B 138B 38.7 26.3 22.1
Together-RPJ 7B 1T 39.4 25.2 21.8
RefinedWeb 7B 138B 39.7 26.5 21.7
LLM360/Amber 7B 1T 39.8 27.9 22.3
FineWeb-Edu 7B 276B 41.9 37.3 24.5
OLMo 7B 2T 42.7 28.4 24.6
Gemma 2B 3T 43.3 40.8 26.6
Falcon 7B 1T 44.1 27.4 25.1
MPT 7B 1T 44.3 28.8 25.7
DCLM-Baseline (ours) 7B 138B 44.8 42.2 0.0
Llama 1 7B 1T 46.3 34.1 28.7
OLMo-1.7 7B 2T 47.0 54.0 34.2
LLM360/CrystalCoder 7B 1T 48.1 48.2 33.2
DCLM-Baseline (ours) 7B 276B 48.7 51.9 32.5
Llama 2 7B 2T 49.2 45.8 34.1
MAP-Neo 7B 4T 50.2 57.1 40.4
DeepSeek 7B 2T 50.7 48.5 35.3
DCLM-Baseline (ours) 7B 3T 57.1 63.7 45.4
Llama 3 8B 15T 57.6 66.2 46.3
Gemma 8B 6T 57.8 64.3 44.6
Appendix R Account of compute costs
@@
We note the compute cost of training runs for each competition scale in Table 1. In total, we estimate that our runs for DCLM sum up to approximately 1.2M H100 hours. Our precise estimate from our experimental test bed is 859K H100 hours for training, but this is likely an underestimate due to additional compute that was not tracked, such as due to training failures. Of the 859K H100 hours, we spent 1,700 on 411M parameter models, 140,000 on 1B models, 4,800 on 3B models, and 713,216 hours on 7B models.
Appendix S Bias and Toxicity Evaluations
We examine the toxicity and bias of DCLM-baseline and compare to other popular base models, running evaluations on several safety-based tasks including CivilComments [27], Copyright [34], BBQ [124], WinoGender [145], and RealToxicityPrompts [65].
We briefly review these benchmarks for those unfamiliar: CivilComments (higher is better) studies how accurately a model can identify toxic content, and we report Exact Match as a metric. Copyright (lower is better) measures the capability of models to reiterate copyrighted content, and we report LCS as a metric. Real Toxicity Prompts (lower is better) measures how easily a user can prompt a model to generate toxic content, and we report Toxic Fraction. BBQ (higher is better) is a question-answering dataset that measures how likely a model’s biases affect its choices, and we report Exact Match. Winogender (higher is better) measures how likely a model is to reinforce a gender-based stereotype when infilling a gendered pronoun.
The results of this analysis can be seen in Table 34. Out datasets lead to models that score comparably to existing ones, trained on preexisting datasets. The model trained on DCLM-Baseline is similar to other popular base models, such as Llama and Mistral, in terms of generating toxic content, as demonstrated by the Real Toxicity Prompts scores. The BBQ and Winogender metrics also demonstrate that our model’s generations tend to contain biases similar to those of other base models. However, our model does not identify toxic comments as accurately as other base models, according to CivilComments. These metrics demonstrate that DCLM-Baseline has relatively similar amounts of bias and toxicity to the datasets that other pretrained models were trained on.
Model CivilComments $\uparrow$ Copyright $\downarrow$ Real Toxicity BBQ $\uparrow$ WinoGender $\uparrow$
Prompts $\downarrow$
DCLM-baseline 0.53 0.01 0.07 0.65 0.62
Llama-2-7B 0.56 0.01 0.09 0.58 0.62
Llama-3-8B 0.74 0.01 0.09 0.67 0.57
Mistral-v0.3-7B 0.67 0.01 0.09 0.71 0.48
Table 34: Bias and toxicity comparisons. We compare the model trained on DCLM-baseline with existing similarly sized base models. We see that, with the exception of CivilComments (toxic comment identification), our model performs similarly to existing ones
Appendix T Existing assets used
@@
In addition to the above, as described in Sections 5, 4.5 and 4.4 and Appendices M and P we make use of the following datasets:
@@
    RedPajama [168] is used for ablations in Section 4.5 and Appendix M. Note that we do not release models or datasets that include this data. RedPajama filters the datasets it uses keeping only permissive licenses, and refers the user to adhere to underlying licenses where appropriate, as described in https://huggingface.co/datasets/togethercomputer/RedPajama-Data-1T.
@@
    • The purpose of DCLM and the associated DCLM-Pool and DCLM-baseline datasets are to enable the study of what makes a strong pretraining dataset for large language models. These models are transformative to society and act as the foundation of numerous applications, but they are often associated with steep costs. While prior work explores many curation techniques, these studies are often coupled with various architectural and training design choices and evaluated in different settings, making controlled comparison nearly impossible. This slows down progress and forces a lot of duplicate work between research teams. Prior work mainly focuses on data curation in the context of supervised datasets and smaller scales (see Section 2 and Appendix B). In our initial release of DCLM, we focus on 53 downstream language understanding tasks that also include reasoning abilities, math, and more. For details see Section 3.5 and Appendix G.
@@
    • DCLM-Pool and DCLM-baseline were created by a group of researchers with the following affiliations, listed in alphabetical order: Allen Institute for Artificial Intelligence, Apple, Bespoke Labs, Carnegie Mellon University, Columbia University, Contextual AI, Cornell University, DatologyAI, Harvard University, Hebrew University, Juelich Supercomputing Center, Research Center Juelich, SambaNova Systems, Stanford University, SynthLabs, Tel Aviv University, Toyota Research Institute, TU Munich, University of California, Los Angeles, University of California, Santa Barbara, University of Southern California, The University of Texas at Austin, University of Washington.
@@
    • We anticipate that DCLM benchmark, tooling and pools will drive data-centric research in ML and AI, fostering the development of the next generation of web-scale datasets, enhancing model abilities, lowering training costs and developing knowledge sharing across research teams.
@@
    • Each instance represents a web-crawled page (document), containing the URL and corresponding HTML content. Each sample is also tagged with metadata about its crawl time and additional information such as the detected language, for processed instances such as those in DCLM-baseline. Additional information can be found in Appendix E.
@@
    • DCLM-Pool contains $\sim$200B documents, all of which are of the same instance type, and comes from hundreds of millions of different sources. The subset DCLM-baseline contains approximately 3B documents.
@@
    • DCLM-Pool is an unfiltered web-text corpus comprised of all Common Crawl data prior to 2023. As such, it represents the full breadth of possible instances from this source. However, we note that Common Crawl does not cover the entire web, due to factors such as compute limitations and opt-outs from page authors. For our DCLM-baseline, we use various filtering and deduplication strategies as described in Section 4 in the explicit attempt to improve its quality for preatining, thus removing low-quality instances, and in doing so, becoming non-representative of the full set of instances. For a complete treatment and visualization of our data processing funnel, see Sections 4, 4.3 and 4.2 and Appendix E.
@@ 8. Q12
    • No. The evaluation procedure is made of tasks as described in Section 3.5. We also attempt to prevent test set contamination in as described in Section 4.6 and Appendix O.
@@ 15. Q19
    • As names and other identifiers are frequent in web data, it is likely that some content can be linked back to specific individuals. However, for most public sites that Common Crawl scrapes, people publish such information willingly, knowing it will be visible and public.
@@ 16. Q20
    • Yes. DCLM-Pool is created from data that is available on the public internet. Since people often debate their political views, sexual preferences, religious beliefs and other such topics, it is highly likely such information is contained in the dataset. While such information is often published willingly with the explicit intent that it will be publicly visible (see Q19), we do encourage additional research on filtering such data both to preserve privacy as well as to discard any potentially biased or toxic content from the training data of downstream models.
@@ 2. Q23
    • We begin by downloading all Common Crawl data prior to 2023. We ran Python-based processing scripts to parse these archives and proprocess them by: (1) filtering low-quality or irrelevant content; (2) deduplicating samples and in some cases decontaminating them against downstream tests sets; (3) computing various model-based features such as fastText classifiers for language detection and quality filtering. We ran processes on hundreds of AWS CPU nodes for Common Crawl parsing and data downloading. Some model-based features (i.e., non fastText models) were computed on GPU clusters. For software links see Q37 and Appendix T or refer to https://datacomp.ai/dclm.
@@ 3. Q24
    • DCLM-Pool is not a probabilistic sample. As described in Q7, DCLM-Pool contains all data from Common Crawl before 2023. Common Crawl is a sample of the Web, and we refer to Common Crawl documentation for details of their sampling process. Our competition pools for the various scales however are fixed subsets from DCLM-Pool, formed by randomly subsetting DCLM-Pool/ Common Crawl paths.